pure acid is to be added to a 20% solution to obtain 60 liters of a 30% acid solution. what amounts of each should be used?
Answer by robertb(5830)

1.00x + 0.20(60-x) = 0.30*60
==> x + 12 - 0.20x = 18
==> 0.80x = 6
==> x = 7.5 liters of pure acid, and
60 - x = 52.5 liters of the 20% acid.